Solve the system of equations and enter as order pairs. Y=-x+5
Y=x-3

Found 3 solutions by Alan3354, stanbon, josh_jordan:
Answer by Alan3354(69443)

Y=-x+5
Y= x-3
----------------- Add

Answer by stanbon(75887)

Solve the system of equations and enter as order pairs.
Y=-x+5
Y=x-3
----------
Substitute for "y" and solve for "x"::
x-3 = -x+5
2x = 8
x = 4
-----
Solve for "y"::
y = x -3
y = 4-3
y = 1
------
Answer:: x = 4 ; y = 1
---------------
Cheers,
Stan H.
--------------

Answer by josh_jordan(263)

1. Rewrite both systems so that the x and y are on the same side of the equal sign and the numbers are on the other side:

-x - y = -5
x - y = 3

2. Add both systems together, since we can easily get rid of x:

-x - y + x - y = -5 + 3 ----->

-2y = -2

3. Divide both sides by -2, giving us our value of y:

y = -2 / -2 = 1

4. Replace the y in either original equation with 1. I'll use the second one:

1 = x - 3

x = 4

So, in ordered pair form, our answer is (4,1).